OpenStudy (anonymous):

The point A(-6,-5) is translated using T: (x,y) -> (x+4,y+6). what is the distance from A to A ?

ganeshie8 (ganeshie8):

change in x = 4 change in y = 6 apply pythagorean theorem : distance = \(\large \sqrt{\Delta x ^2 + \Delta y^2} =\sqrt{4^2 + 6^2} = ?\)

OpenStudy (anonymous):

7.2 ?

ganeshie8 (ganeshie8):

Yep!